Hi please help on doing these type of equations :)
Question 1
6b + 4y = 54
4b = 20
Question 2
2b - 5x = -19
-b= -3
Question 3
3u - 3y = 30
5u= 20
Thank you

Answer by ohtheirony(35)

To solve these problems, you can use substitution method. Take the small equation and put it into the large equation.


Problem 1: solve the 5b=20, b=4
Now, if we put b=20 into the 6b+4y=54, we can solve for y.
6(20)+4y=54
120+4y=54
4y=-66
y=-16.5


Problem 2: -b=-3 becomes b=3
2(3)-5x=-19
6-5x=-19
-5x=-25
x=5


Problem 3: 5u=20 becomes u=4
3(4)-3y=30
12-3y=30
-3y=18
y=-6


When you get to longer equations, it may be helpful to look at this tutorial:
http://www.basic-mathematics.com/substitution-method.html
But don't forget about elimination! That can work just as well for system of equations.
http://www.purplemath.com/modules/systlin5.htm